An RLC series circuit has R = 100 ohms, C = 25 µF, and L = 0.16 H. For what angular
frequency of an ac voltage is the current flow maximum?
1
Expert's answer
2020-11-06T11:27:03-0500

In an RLC series circuit for maximum current to flow through the circuit. The capacitive reactance"(X_C)" should be equal to the inductive reactance "(X_L)" . This condition is called the Resonance.


Therefore we have,

"X_L = X_C"

"L\omega = \frac{1}{C\omega}"


"\space\space\space\space \space\space \omega^2 = \frac{1}{LC}"


"\omega= \frac{1}{25\times10^{-6}\times0.16}"


"\omega=500\space rad/sec"